how can we calculate the rate of effusion of nitrogen gas from a cylinder. and how much time it will take??

Grade:11

3 Answers

Akash Kumar Dutta
11 years ago

Dear Rajveer,
Apply Grahams law of effusion.
it is:-
rate(r) of effusion/diffusion is proportional to 1/ root [density(d) ]
hecne r1/r2 = root(  d2/d1 )
         => r1/r2 = root ( M2/M1 )   since   [2. vapour density(d)=molecular mass(M) ]

        now put rate=Volume/time=V/t
        all sums of this type are solved by this concept.
Hope it gets cleared now.

Regards.

Prajwal kr
11 years ago

r=(volume of gas)/time taken.

ANUBHAV SINHA
11 years ago

rate of effusion=r=(volume of nitrogen gas effusing from the cylinder)/(time taken)

also since pressure and temperature are constant, volume is proprtional to no. of moles

 

so,r=(no. of moles of N2 effusing)/(time taken)